# Diffusion Time-step Curriculum for One Image to 3D Generation

### v2

@@ Abstract
footnotetext: $*{}^{*}$orresponding author.
@@ 1 Introduction
  • We develop an end-end one image-3D pipeline DTC123, boosting the efficiency, diversity and fidelity of prevalent SDS-based methods in both real-world and synthetic scenarios.
@@ 2 Related Work
Image-3D Generation focuses on generating 3D assets from a given reference image, which can also be considered as an ill-posed single-view reconstruction problem [7]. The above text-to-3D methods can be directly adapted for image-to-3D generation with image captioning, e.g., Realfusion [35] and NeRDi [7] directly adapted SDS [39] into Image-to-3D with textual inversion of the given image. Magic123 [40] and Consistent123 [61] combine Zero-1-to-3 [30] with RealFusion [35] to further improve the generation quality. One-2-3-45 [30] and Dreamgaussian [54] adopted NeuS [58, 59] and Generative Gaussian Splatting [23] as the 3D representation respectively, significantly reducing the generation time at the expense of the generation quality. We build upon these optimization-based approaches by implementing a coarse-to-fine optimization strategy that explores the diffusion time-step curriculum they overlooked, which enables efficient generation of high-fidelity, multi-view consistent 3D assets from a single image.
@@ 3 Preliminary
Given a single image of an object, our goal is to optimize a coherent 3D model (e.g., NeRF, mesh) so that it can restore the given image from the reference view $\pi_{\text{ref}}$ and generate a highly plausible image from any unseen view $\pi$ with the supervision of the diffusion model. Here, we mainly introduce Score Distillation Sampling (SDS)  [39] with student 3D models, teacher diffusion models (see Appendix C) that will help to build up our approach.
@@ 3.1 Score Distillation Sampling (SDS)
$\small\nabla_{\bm{\theta}}\mathcal{L}_{\text{SDS}}(\bm{\theta},t)=$ $\mathbb{E}_{\bm{\epsilon}}\left[\omega(t)\left(\bm{\epsilon}_{\phi}\left(\mathbf{x}_{t};t,\bm{y}\right)-\bm{\epsilon}\right)\frac{\partial\mathbf{x}}{\partial\bm{\theta}}\right]$
@@
where $\small\hat{\mathbf{x}}_{0}=\mathbf{x}_{t}-\sigma_{t}\bm{\epsilon}_{\phi}\left(\mathbf{x}_{t};\bm{y},t\right)$, can be considered as single-step de-nosing output with starting point $\mathbf{x}_{t}$; $\bm{y}$ is the condition (e.g., text, camera pose) and depends on the types of teacher diffusion models; $\bar{{\omega}}(t)=\omega(t)/\sigma_{t}$, denotes the weight function. Thus, we reveal that the crux of this teacher-student optimization process is directly determined by such perturbed-and-denoised output, i.e., the quality of the teacher-generated $\hat{\mathbf{x}}_{0}$. As illustrated in Figure 2(a), intuitively, not all time-steps $t$ can provide useful and valid guidance $\hat{\mathbf{x}}_{0}$, which motivates the following proposed diffusion time-step curriculum.
@@ 4 DTC123
In this section, we introduce our diffusion time-step curriculum one-image-to-3D pipeline, called “DTC123”, with a theoretical justification in Appendix A. As shown in Figure 3(a): We start by taking a reference image and extracting its geometric estimation and text description. Then the optimization procedure can be categorized into unseen view guidance and reference restoration:
@@ 4.1 Annealed Time-Step Schedule
We develop a time-step annealed sampling schedule for our time-step curriculum, thereby facilitating teacher-student knowledge transfer in a progressive, coarse-to-fine manner. Please refers to Appendix A for its theoretical justification. In particular, we use an annealed interval $[t_{\text{mid}}-\Delta,t_{\text{mid}}+\Delta]$ to randomly select the time-step $t$, with the interval midpoint $t_{\text{mid}}$ decreasing monotonically. The interval radius $\Delta$ narrows progressively during optimization, and the midpoint $t_{\text{mid}}$ is determined by the current iteration $k$:
@@
where $t_{\max}$ and $t_{\min}$ refers to the entire diffusion time-step range; $l$ and $N$ denotes the step length and total training iteration, respectively. This schedule not only introduces local randomness to preserve the vibrancy of the model’s coloration but also allocates more iterations to smaller time-steps for in-depth detail exploration. Empirical observation in Appendix B demonstrates the superiority and robustness of our sampling schedule.
@@ 4.3 Teacher: Coarse-to-fine Diffusion Prior
The teacher diffusion model should follow the diffusion time-step curriculum by initially offering a rough silhouette of the object in the desired pose in larger time steps, subsequently prioritizing intricate texture refinement in small time steps. A natural question follows : What is the suitable teacher for the time-step curriculum? Empirical evidence [^1] To answer this question, we quantitatively analyze two most popular teacher diffusion models by investigating the contour exploration consistency via MaskIoU and the perceptual generation quality computed by CLIP-R [29] in the OmniObject3D [64] dataset. See detailed experimental results in Appendix B. suggests: the view-conditioned Zero-1-to-3 serves as a coarse-grained teacher by providing a more accurate geometry structure at large time-step. Conversely, the text-conditioned Stable Diffusion is suitable for a fine-grained teacher as it yields realistic texture details at smaller time-step.
@@
where the time-step $t$ follows the sampling schedule in Sec. 4.1; $\lambda$ is the trade-off hyper-parameter, set as 0 in the first stage and gradually increase with the reducing time-step in the second stage for finer detail guidance. We apply DDIM sampling [49, 9] for multi-step de-nosing process, i.e., $\mathbf{x}_{t}\rightarrow\mathbf{x}_{t-r}\ldots\rightarrow\hat{\mathbf{x}}_{0}$ instead of the imprecise single-step de-noising in the last few iterations for both efficiency and performance concerns. Empirically, it equips the teacher-generated $\hat{\mathbf{x}}_{0}$ with more decent details and mitigates issues like texture flickering or overly saturated color blocks at smaller time-steps. In addition, we propose the following approaches to further improve the fine-grained supervision quality in scenarios with sophisticated textures by clarifying the ambiguous inferred text description and alleviating the Janus face problem.
@@ 5.2 Experimental Protocol
An exceptional 3D model should not only mirror the reference image, but also maintain a consistent correlation with the reference and plausible results when observed from alternative poses. Following  [40], we compared our DTC123 in RealFusion15 [35], NeRF4 benchmark for quantitative comparison, with PSNR and LPIPS metrics to measure reconstruction quality and CLIP-similarity to evaluate appearance similarity. For qualitative results, we manually collected 50 reference images from the Internet out of the range of Objaverse, covering a wider range of items with various difficulty.
We adopted Zero123-SDS [30], RealFusion [35], NeuralLift [68] and Magic123 [40] as baseline SDS-based methods with their default experimental settings. We also compared with the state-of-the-art methods One-2-3-45 [29] and SyncDreamer [32], which perform Image-3D generation in a feed-forward manner instead of SDS optimization. For Zero123-SDS and Magic123, we adopted the implementation from threestudio [14] and leveraged Zero123-XL as ours for fair comparison, while others from their official codebase. We are quite confident that the baselines presented here are the finest re-implementations we have come across. Note that our proposed diffusion time-step curriculum could also be adapted to text-to-3D tasks by replacing Zero-1-to-3 with MVdream for coarse guidance and leveraging different time-step schedule and prompt de-biasing [17].
@@ 5.3 Image-3D Generation
Qualitative Results. Figure 5 demonstrates that DTC123 maintains high fidelity and plausible generation in complex scenarios. In contrast, generation results from most baselines, even those utilizing the more advanced Zero123-XL, are plagued by multi-view inconsistency, geometric distortion, and texture conflict. For example, in the yellow robot case in the fourth line of Figure 5, competing textures cause blurring and exhibit unreasonable features, such as an extra eye on the back. In comparison, DTC123 generates high-fidelity novel views with realistic rendered metallic textures.
@@
Table 1: Quantitative results. We show quantitative results in terms of CLIP-Similaritynormal-$\uparrow$ / PSNRnormal-$\uparrow$ / LPIPSnormal-$\downarrow$. The results are shown on the NeRF4 and RealFusion datasets.
@@ 6 Ablation and Discussion
A2: As depicted in Figure 7, DTC123 consistently exhibits a lower failure rate compared to other methods across various initialization (random seeds) and at different difficulty levels by a large margin. Notably, DTC123 maintains a degree of robustness even at the most challenging Level 5, characterized by complex illumination and multi-object topology. Such robustness should be attributed to the proposed time-step curriculum, where the student 3D model initially captures coarse features, significantly reducing the instability typically associated with randomized NeRF initialization.
@@
A3: We clarify that annealing time-step is NOT equivalent to, but a part of time-step curriculum. Our DTC includes three coherent parts : time-step schedule, progressive student representation and teacher guidance. Different from the viewpoint of DDPM sampling from Dreamtime [18], we consider that SDS leverages teacher-generated $\hat{\mathbf{x}}_{0}$, which is a single-step de-nosing output with starting point $\mathbf{x}_{t}$, to optimize student-rendered $\mathbf{x}_{\pi}$. Modeling 3D generation as a data corruption reduction process, we reveal that there exists a time-step lower bound in SDS for teacher diffusion model to well estimate the desired score function and thus provide quality guidance $\hat{\mathbf{x}}_{0}$. In order to alleviate the marked divergence of the teacher-generated $\hat{\mathbf{x}}_{0}$ [16] at large time-steps with such schedule, we further design the progressive student representation and teacher guidance to cooperate with the annealing time-step for a more stable coarse-to-fine generation. Please refers to Appendix A for more details.
Q4: Is there any limitations of DTC123?
A4: Though we are quite confident that our proposed DTC123 shows consistent superiority over other methods, it is important to acknowledge several remaining limitations: (1) Teacher diffusion models may become a bottleneck, introducing unwanted bias [71, 17] in certain scenarios. For instance, if the Zero-1-to-3 model fails to accurately capture the basic structure corresponding to the desired pose in the first stage, the results are likely to be unsatisfactory. (2) Our approach presumes that the reference image is from a front view; deviations from this assumption can lead to significant failures in the generation results. We propose that integrating vision-language models (VLMs) or employing manual camera estimation could potentially address this limitation.
@@ 7 Conclusion
We revisit Score Distillation Sampling (SDS) and point out that the crux of its enhancement lies in the proposed diffusion time-step curriculum. We then design an improved coarse-to-fine Image-3D pipeline (DTC123) which collaborates the teacher diffusion model and 3D student model with the time-step curriculum. Through extensive qualitative comparisons and quantitative evaluations, we show that our DTC123 significantly improves the photo-realism and multi-view consistency of Image-to-3D generation. In the future, we will focus on exploring the potential of diffusion time-step curriculum with advanced teacher models and diverse student models to further improve the generation efficiency and quality.
@@ Appendix A Theoretical Justification
[…]

### v3

@@ Abstract
footnotetext: Corresponding author.
@@ 1 Introduction
  • We develop an end-end one image-3D pipeline DTC123, boosting the efficiency, diversity and fidelity of SDS-based methods in real-world and synthetic scenarios.
@@ 2 Related Work
Image-3D Generation focuses on generating 3D assets from a given reference image, which can also be considered as an ill-posed single-view reconstruction problem [7]. The above text-to-3D methods can be directly adapted for image-to-3D generation with image captioning, e.g., Realfusion [35] and NeRDi [7] directly adapted SDS [39] into Image-to-3D with textual inversion of the given image. Magic123 [40] and Consistent123 [61] combine Zero-1-to-3 [30] with RealFusion [35] to further improve the generation quality. One-2-3-45 [30] and Dreamgaussian [54] adopted NeuS [58, 59] and Generative Gaussian Splatting [23] as the 3D representation respectively, significantly reducing the generation time at the expense of the generation quality. We build upon these optimization-based approaches by implementing a coarse-to-fine optimization strategy that explores the diffusion time-step curriculum they overlooked, which enables efficient generation of high-fidelity, multi-view consistent 3D assets from one image.
@@ 3 Preliminary
Given a single image of an object, our goal is to optimize a coherent 3D model (e.g., NeRF, mesh) so that it can restore the given image from the reference view $\pi_{\text{ref}}$ and generate a highly plausible image from any unseen view $\pi$ with the supervision of the diffusion model. Here, we mainly introduce Score Distillation Sampling (SDS)  [39] with student 3D models, teacher diffusion models (see Appendix) that will help to build up our approach.
@@ 3.1 Score Distillation Sampling (SDS)
$\nabla_{\bm{\theta}}\mathcal{L}_{\text{SDS}}(\bm{\theta},t)=$ $\mathbb{E}_{\bm{\epsilon}}\left[\omega(t)\left(\bm{\epsilon}_{\phi}\left(\mathbf{x}_{t};t,\bm{y}\right)-\bm{\epsilon}\right)\frac{\partial\mathbf{x}}{\partial\bm{\theta}}\right]$
@@
where $\hat{\mathbf{x}}_{0}=\mathbf{x}_{t}-\sigma_{t}\bm{\epsilon}_{\phi}\left(\mathbf{x}_{t};\bm{y},t\right)$, can be considered as single-step de-nosing output with starting point $\mathbf{x}_{t}$; $\bm{y}$ is the condition (e.g., text, camera pose) and depends on the types of teacher diffusion models; $\bar{{\omega}}(t)=\omega(t)/\sigma_{t}$, denotes the weight function. Thus, we reveal that the crux of this teacher-student optimization process is directly determined by such perturbed-and-denoised output, i.e., the quality of the teacher-generated $\hat{\mathbf{x}}_{0}$. As illustrated in Figure 2(a), intuitively, not all time-steps $t$ can provide useful and valid guidance $\hat{\mathbf{x}}_{0}$, which motivates the following proposed diffusion time-step curriculum.
@@ 4 DTC123
In this section, we introduce our diffusion time-step curriculum one-image-to-3D pipeline, called “DTC123”, with a theoretical justification in Appendix. As shown in Figure 3(a): We start by taking a reference image and extracting its geometric estimation and text description. Then the optimization procedure can be categorized into unseen view guidance and reference restoration:
@@ 4.1 Annealed Time-Step Schedule
We develop a time-step annealed sampling schedule for our time-step curriculum, thereby facilitating teacher-student knowledge transfer in a progressive, coarse-to-fine manner. Please refers to Appendix for its theoretical justification. In particular, we use an annealed interval $[t_{\text{mid}}-\Delta,t_{\text{mid}}+\Delta]$ to randomly select the time-step $t$, with the interval midpoint $t_{\text{mid}}$ decreasing monotonically. The interval radius $\Delta$ narrows progressively during optimization, and the midpoint $t_{\text{mid}}$ is determined by the current iteration $k$:
@@
where $t_{\max}$ and $t_{\min}$ refers to the entire diffusion time-step range; $l$ and $N$ denotes the step length and total training iteration, respectively. This schedule not only introduces local randomness to preserve the vibrancy of the model’s coloration but also allocates more iterations to smaller time-steps for in-depth detail exploration.
@@ 4.3 Teacher: Coarse-to-fine Diffusion Prior
The teacher diffusion model should follow the diffusion time-step curriculum by initially offering a rough silhouette of the object in the desired pose in larger time steps, subsequently prioritizing intricate texture refinement in small time steps. A natural question follows : What is the suitable teacher for the time-step curriculum? Empirical evidence [^1] To answer this question, we quantitatively analyze two most popular teacher diffusion models by investigating the contour exploration consistency via MaskIoU and the perceptual generation quality computed by CLIP-R [29] in the OmniObject3D [64] dataset. See detailed experimental results in Appendix. suggests: the view-conditioned Zero-1-to-3 serves as a coarse-grained teacher by providing a more accurate geometry structure at large time-step. Conversely, the Stable Diffusion is suitable for a fine-grained teacher as it yields realistic texture details at smaller time-step.
@@
where the time-step $t$ follows the sampling schedule in Sec. 4.1; $\lambda$ is the trade-off hyper-parameter, set as 0 in the first stage and gradually increase with the reducing time-step in the second stage. We apply DDIM sampling [49, 9] for multi-step de-nosing process, i.e., $\mathbf{x}_{t}\rightarrow\mathbf{x}_{t-r}\ldots\rightarrow\hat{\mathbf{x}}_{0}$ instead of the imprecise single-step de-noising in the last few iterations for both efficiency and performance concerns. Empirically, it equips the teacher-generated $\hat{\mathbf{x}}_{0}$ with more decent details and mitigates issues like texture flickering or overly saturated color blocks at smaller time-steps. In addition, we propose the following approaches to further improve the fine-grained supervision quality in scenarios with sophisticated textures by clarifying the ambiguous inferred text description and alleviating the Janus face problem.
@@ 5.2 Experimental Protocol
An exceptional 3D model should not only mirror the reference image but also maintain a consistent correlation with the reference and plausible results when observed from other poses. Following  [40], we compared our DTC123 in RealFusion15 [35], NeRF4 benchmark for quantitative comparison, with PSNR and LPIPS metrics to measure reconstruction quality and CLIP-similarity to evaluate appearance similarity. For qualitative results, we manually collected 50 reference images from the Internet out of the range of Objaverse, covering a wider range of difficult items.
We adopted Zero123-SDS [30], RealFusion [35], NeuralLift [68] and Magic123 [40] as baseline SDS-based methods with their default experimental settings. We also compared with the state-of-the-art methods One-2-3-45 [29] and SyncDreamer [32], which perform Image-3D generation in a feed-forward manner instead of SDS optimization. For Zero123-SDS and Magic123, we adopted the implementation from threestudio [14] and leveraged Zero123-XL as ours for fair comparison, while others from their official codebase. We are quite confident that the baselines presented here are the finest re-implementations we have come across.
@@ 5.3 Image-3D Generation
Qualitative Results. Figure 5 demonstrates that DTC123 maintains high fidelity and plausible generation in complex scenarios. In contrast, generation results from most baselines, even those utilizing the more advanced Zero123-XL, are plagued by multi-view inconsistency, geometric distortion, and texture conflict. For example, in the yellow robot case in the fourth line of Figure 5, competing textures cause blurring and exhibit unreasonable features, such as an extra eye on the back. In comparison, DTC123 generates high-fidelity novel views with realistic metallic textures.
@@
[…]
@@ 6 Ablation and Discussion
A2: As depicted in Figure 7, DTC123 consistently exhibits a lower failure rate compared to other methods across various initialization (random seeds) and at different difficulty levels by a large margin. Such robustness should be attributed to the proposed time-step curriculum, where the student 3D model initially captures coarse features, significantly reducing the instability typically associated with randomized NeRF initialization.
@@
A3: We clarify that annealing time-step is NOT equivalent to, but a part of time-step curriculum. Our DTC includes three coherent parts : time-step schedule, progressive student representation and teacher guidance. Different from the viewpoint of DDPM sampling from Dreamtime [18], we consider that SDS leverages teacher-generated $\hat{\mathbf{x}}_{0}$, which is a single-step de-nosing output with starting point $\mathbf{x}_{t}$, to optimize student-rendered $\mathbf{x}_{\pi}$. Modeling 3D generation as a data corruption reduction process, we reveal that there exists a time-step lower bound in SDS for teacher diffusion model to well estimate the desired score function and thus provide quality guidance $\hat{\mathbf{x}}_{0}$. In order to alleviate the marked divergence of the teacher-generated $\hat{\mathbf{x}}_{0}$ [16] at large time-steps, we further design the progressive student representation and teacher guidance to cooperate with the annealing time-step for a more stable coarse-to-fine generation. Please refers to Appendix for more details.
@@ 7 Conclusion
We revisit Score Distillation Sampling (SDS) and point out that the crux of its enhancement lies in the proposed diffusion time-step curriculum. We then design an improved coarse-to-fine Image-3D pipeline (DTC123) which collaborates the teacher diffusion model and 3D student model with the time-step curriculum. Through qualitative comparisons and quantitative evaluations, we show that our DTC123 significantly improves the photo-realism and multi-view consistency of Image-to-3D generation. In the future, we will focus on exploring the potential of diffusion time-step curriculum with advanced teacher models and diverse student models to further improve the efficiency and quality.
@@ Appendix A Theoretical Justification
(Data Corruption Reduction) Given the camera pose $\pi$ and the condition $y$, we consider a student-rendered image $\mathbf{x}_{\pi}=g(\theta,\pi)$ at an arbitrary training iteration $k$. Suppose that there exists a real data point $\mathbf{x}^{*}$ drawn from the data distribution $p_{\text{data}}(\mathbf{x})$ and an unknown data corruption $\delta_{k}=D_{\mathrm{KL}}[\delta(\mathbf{x}-\mathbf{x}_{\pi})||p_{\text{data}}(\mathbf{x})]$, such that we can express $\mathbf{x}_{\pi}$ as $\mathbf{x}_{\pi}=\mathbf{x}^{*}+\delta_{k}$. Our objective is to iteratively reduce the corruption $\delta_{k}$ inherent in $\mathbf{x}_{\pi}$ as $k\to\infty$, guided by the conditional score function $\nabla\log p_{t}(\mathbf{x})$, such that $\mathbf{x}_{\pi}$ increasingly resembles a sample from $p_{\text{data}}(\mathbf{x})$.